zarkam21:

velocity of a neutron with 100pm wavelength

zarkam21:

@Vocaloid

Vocaloid:



Vocaloid:

make sure to convert pm to m first

zarkam21:

but what would the mass be?

Vocaloid:

mass of a neutron in kg

zarkam21:

okay so i woud first rearrange the equation to v=h/(m)(lambda)

Vocaloid:

yes

zarkam21:

is the rearragement good

Vocaloid:

yes

zarkam21:

3.956x10^-21

Vocaloid:

as your final solution? hm, no, a neutron would move much faster than that 100pm * (10^(-12) will convert pm to meters so 100 * 10^(-12) = (6.626e-34)/(v * 1.674927351e-27) solve for v

zarkam21:

3656

Vocaloid:

hm, I got something slightly different chucking it into a calculator gives me 3955.99 m/s (they only gave you 1 sig fig so round up to 4000?)

zarkam21:

Yeah I would believe so

zarkam21:

lightbulb can produce 9.8x10^20 photons in one second with a wavelength of 434nm. Hoe much energy in joules will this light build produce in one hour

zarkam21:

okay so first is to convert nm to m and I got 10x10^-18

zarkam21:

and the formula would be E=hv=hc/lambda

Vocaloid:

hm. not 100% sure but here are my thoughts: E = hc/lambda plugging in planck's constant, wavelength, speed of light to get E for one photon then multiply by the number of photons 9.8x10^20 photons to get joules then, since that's energy per second, multiply by 3600 to get energy per hour since 1 hour = 3600 seconds

zarkam21:

wave length would be 10x10^-18

Vocaloid:

hm not quite 1 nm = 10*(-9) m so 434 nm = 434 * 10^(-9) m

zarkam21:

1.615x10^-8 energy/second

Vocaloid:

hm, not quite, you may have inverted something anyway E = hc/lambda = 6.626e-34*3e8/(434*10^(-9)) multiplying by the number of photons, then multiplying by 3600 6.626e-34*3e8/(434*10^(-9)) * 3600 * 9.8x10^20

zarkam21:

for some reason I'm getting 0 x 10^20

Vocaloid:

hm. might be an exponent issue w/ whatever program you're using anyway I end up with 1.6e6

zarkam21:

Joules

Vocaloid:

yes

zarkam21:

spectral lines in the Balmer series, emission back down to the n=2 level, for the hydrogen atom occur at 656.3nm, 486.1nm, 434.0nm, and 410.2nm. What is the wavelength for the next line in this series

Vocaloid:

hm. i'd have to think about this one for a bit. you'd probably use the rydberg equation where nf is 2 at least.

zarkam21:

this would be the En=rhc/n^2?

Vocaloid:

hm, close but not quite. we are dealing with wavelengths not energies so 1/lambda = R(1/nf^2 - 1/ni^2) anyway I got what they were going for. if you let nf = 2, and plug in ni = 3,4,5, and 6, you end up with the wavelengths they give you so the next one must be ni = 7 1/lambda = (1.097e-7)(1/2^2 - 1/7^2) solve for lambda

zarkam21:

3.x10^-9

zarkam21:

3.9x10^-9

Vocaloid:

hm, I got something a bit different, (397nm) i'm using wolframalpha for my calculations

zarkam21:

do i vhange it to m

Vocaloid:

there's an unspoken rule in math that you should give back whatever units they give you so since they give you nm in the original equation stick with nm